OpenStudy (anonymous):

What is the equation of the line that is parallel to the line y = 2x + 2 and passes through the point (5, 3)?

OpenStudy (anonymous):

okay pararlle y=2x+b 3=2*5+b -7=b so the eqn y=2x-7

OpenStudy (anonymous):

thank you!!